OpenStudy (anonymous):

Which of the following equations represents an inverse variation? A. B. C. D.

OpenStudy (anonymous):

a.xy=5 b.5y=x y/x=5 x/y=5

OpenStudy (luigi0210):

Inverse variations usually include K?

OpenStudy (anonymous):

nope i typed it like it was @Luigi0210

OpenStudy (luigi0210):

Well, USUALLY it's something like \[\LARGE y=\frac{k}{x}\] or something. I guess it just needs to have some sort of constant included and y/x

OpenStudy (anonymous):

so non e of these 4 will work? @Luigi0210

OpenStudy (luigi0210):

Actually, if you multiply it by x we get.. \[\LARGE xy=k\] Where k is any constant so it could 5